### v1

# Meta-Property Graphs: Extending Property Graphs with Metadata Awareness and Reification

Abstract
The ISO standard Property Graph model has become increasingly popular for representing complex, interconnected data. However, it lacks native support for querying metadata and reification, which limits its abilities to deal with the demands of modern applications. We introduce the vision of Meta-Property Graph, a backwards compatible extension of the property graph model addressing these limitations. Our approach enables first-class treatment of labels and properties as queryable objects and supports reification of substructures in a graph. We propose MetaGPML, a backwards compatible extension of the Graph Pattern Matching Language forming the core of the ISO standard GQL, to query these enhanced graphs. We demonstrate how these foundations pave the way for advanced data analytics and governance tasks that are challenging or impossible with current property graph systems.
1 Introduction
Motivation.
Modern data engineering applications demand flexibility and agility more than ever before. For example, during exploratory analytics, one typically starts in a schema-less situation, where structure is discovered over time. Heterogeneity in what is a data value versus what is an attribute name is inherent during discovery, profiling, and exploration. As another example, during integration of data sources, such as when building a knowledge graph, data and schema heterogeneity is the default. One source’s data values correspond to another source’s node labels, and in yet another data source, these correspond to node property names; further, what is a subgraph in one source corresponds to a node in a different data source. Modern data management solutions must fully embrace and support heterogeneity and diversity in the modeling distinctions between data and metadata.
In data modelling, metadata is commonly understood in two distinct ways. The first and most common form is attribute data, which refers to the characteristics or details associated with a piece of information, such as a relational database table’s attributes (i.e., column names) or the table’s name itself. For instance, when storing a person’s contact information, attributes like name, email address, phone number, and address are essential metadata. On the other hand, another common form of metadata is reification. This concept involves aggregating complex relationships or sets into a new entity or object, making it easier to manage and analyze. In the context of entity-relationship modeling, reification is often used to transform relationship sets into entity sets, allowing for more effective data modeling and manipulation. For example, in an online shopping platform, reifying the relationship between customers and orders could result in a new entity representing the order history or customer preferences. [^1] Other more complex types of metadata, such as reflective or active data [7, 9, 18], are beyond the scope of our current discussion.
The ISO standard Property Graph (PGs) model has gained popularity in graph data management and is widely adopted, e.g., in graph DB systems such as Neo4j, Tigergraph, and Amazon Neptune, as well as relational DB systems such as DuckDB which implement the ISO extensions to SQL for PG querying. In a property graph nodes and edges are labeled and also have associated sets of property name/value pairs (e.g., a node labeled Person with property Birthdate having data value 11-11-2001; here Person and Birthdate are attribute metadata). While PGs offer a model closely aligned with conceptual domain representations, the model makes (1) a strict distinction between metadata and data, and (2) has no support for reification. In contrast, the W3C’s Resource Description Framework (RDF) for modeling graph data treats everything as data, including attributes of nodes and edges, and natively supports reification. This capability makes RDF particularly useful for applications requiring sophisticated data validation, verification, querying, and governance of metadata. The lack of metadata awareness in PGs, such as the ability to query and audit metadata, represents a significant limitation compared to RDF’s versatility.
The vision.
In this paper we present a vision for overcoming barriers to flexible management of data-metadata heterogeneity in PG data management applications. Towards this, we introduce Meta-Property Graphs (MPG), a fully backwards compatible extension of the PG model that addresses limitations in representing and querying metadata. Our approach enables first-class treatment of labels and properties as queryable objects, as well as reification of subgraphs. On this foundation, we further propose MetaGPML, an fully backwards compatible extension of the Graph Pattern Matching Language (GPML), the core language at the heart of the ISO standard GQL for PG querying. We give a complete formal specification of MPG and MetaGPML, providing the foundations for our vision. Furthermore, we demonstrate how these contributions facilitate advanced data analytics, integration, and governance tasks that are challenging or impossible with current PG systems.
2 Related Work
The PG model represents a design point on the continuum between the Relational (fixed metadata) and RDF (no fixed metadata) data models. While by default schema-less, supporting agile modeling and development, PGs nonetheless make a strict distinction between data values and metadata values. Of course, all points along this design continuum are equally important, finding their applications and use-cases. And indeed, data cleaning, wrangling, integration, and exchange are very often about moving (meta)data along the continuum. What has been missing is an appropriate design for PGs to fully meet modeling and querying needs of data management scenarios such as these (see Chapter 2 of [5] for a survey of design approaches for PG modeling).
In the research literature, dealing with the challenges of data-metadata heterogeneity was studied in the context of relational data integration and data integration on the web, leading to solutions for relational and XML data-metadata mapping and exchange (e.g., [4, 10, 15]) and seminal work such as the SchemaSQL and FISQL data-metadata query languages for relational databases [16, 21]. Our work approaches these challenges in the new context of the ISO standards for graph data management.
In another direction, there is a growing body of work on mappings between graph data models [1, 6, 14, 20, 19] and generalizing both PG and RDF [13, 17, 3]. Our vision is orthogonal and complementary to these investigations, aiming at extending the PG model and the standardized GQL and SQL/PGQ query languages with seamless data-metadata functionalities.
3 A guided tour of working with MPGs
Figure 1: Example meta-property graph $G$
Figure 1 illustrates a simple MPG, demonstrating how the MPG model enables direct querying of metadata as data objects. The sample database represents publications, indexing databases, and persons with their relationships, modeled using four types of data objects: edges ($E_{G}$), nodes ($N_{G}$), properties ($P_{G}$), and labels ($L_{G}$). Note that properties and labels are now first class citizens, unlike in the current PG standard.
3.1 Working with Data, Labels, and Properties
Meta-Property Graph enables us to query directly on label sets as recognized data objects without considering the nodes or edges that they are associated with. This will facilitate analytics over classes in the database and work with this data independently. As a simple instance, in the following example MetaGPML query $Q_{1}$, we retrieve the label set that contains labels Publication in the database to figure out the other co-occurring tags regardless of which node they belong to.
$Q_{1}$: Which label sets contain the label Publication? Result table of $Q_{1}$
Publication_Co_Tags
{”Publication”, ”Journal”}
{”Publication”, ”Conference”}
Here l is a variable which binds to label set objects. With MetaGPML we can also query properties as independent data objects. Query $Q_{2}$ demonstrates this by retrieving the values of all Name properties, regardless of whether the property is attached to a node or an edge. $Q_{2}$: What are the values of Name properties? Result table of $Q_{2}$
Names
Lee
Scopus
Rose
PubMed
[…]
3.2 Working with Reification
MPG and MetaGPML enable reification as the second type of metadata on property graphs. A sub-structure of an MPG can be reified as a node, thereby making it a first class citizen. This flexibility allows for partial relationship reification, which can be useful for privacy preservation. The model also permits a node to reify its own properties alongside other data objects. In Figure 1, we represent the statement “Rose assigned Lee as a reviewer on 5th November 2024” using a node that induces a sub-structure containing node $n_{1}$, relevant properties, and outgoing review edges. Query $Q_{5}$ retrieves the name of the person who assigns a reviewer:
$Q_{5}$: Who assigned Lee as a reviewer and when? Note that in the MATCH clause we have a graph pattern embedded in a node pattern, to denote a query to be executed on the sub-structure reified by the node which is bound to variable y. Result table of $Q_{5}$
Reviewer name Date Assigning editor
Lee  05-11-2024 Rose
4 Meta-Property Graph Model
We next formalize Meta-Property Graphs. Let $\mathcal{I}$, $\mathcal{L}$, $\mathcal{K}$, and $\mathcal{V}$ be pairwise disjoint sets of object identifiers, labels, property keys, and property values, respectively.
Definition 4.1
A meta-property graph is a directed and undirected vertex- and edge-labeled graph $G=(N,E,P,L,\lambda,\mu,\sigma,\upsilon,\eta,\rho)$, where:
  • $N,E,P,L\subseteq\mathcal{I}$ are finite, pairwise disjoint sets,
  • $\lambda:N\cup E\rightarrow L$ is a bijective labeling function,
  • $\mu:L\rightarrow 2^{\mathcal{L}}$ assigns a finite set of labels to each label identifier,
  • $\upsilon:P\rightarrow\mathcal{K}\times\mathcal{V}$ assigns key-value pairs to properties,
  • $\sigma:N\cup E\rightarrow\mathbb{C}$ assigns compatible property sets to nodes and edges, such that for each pair of distinct objects $o_{1},o_{2}\in N\cup E$, it holds that $\sigma{(o_{1})}\cap\sigma{(o_{2})}=\varnothing$ and $\bigcup_{o\in N\cup E}\sigma(o)=P$, i.e., every and each property $p\in P$ is assigned to exactly one node or edge.
  • $\eta=(\eta_{s},\eta_{t},\eta_{u})$ where:
    • $\eta_{s},\eta_{t}:E_{d}\rightarrow N$ assign source and target nodes to directed edges,
    • $\eta_{u}:E_{u}\rightarrow\{\{u,v\}\mid u,v\in N\}$ assigns node pairs to undirected edges,
  • $\rho:N\rightarrow 2^{N\cup E\cup P\cup L}$ assigns finite sets of objects to nodes, such that for each $n\in N$, it holds that $n\notin\rho^{*}(n)$, where $\rho^{*}(n)$ is the closure of $\rho(n)$, [^2] Formally, $\rho^{0}(n)=\rho(n)$, $\rho^{i}(n)=\rho^{i-1}(n)\cup\bigcup_{n^{\prime}\in N\cap\rho^{i-1}(n)}\rho(n^{\prime})$, and $\rho^{*}(n)=\bigcup_{i=0}^{\infty}\rho^{i}(n)$. ensuring that the sub-structure associated with each node is well-founded.
where $E=E_{d}\cup E_{u}$ the set of compatible property sets =$\mathbb{C}=\{C\subseteq P\mid\forall p_{1},p_{2}\in C,p_{1}\neq p_{2}\Rightarrow Key(p_{1})\neq Key(p_{2})\}$, $Key(p)=\pi_{1}(\upsilon(p))$, $Val(p)=\pi_{2}(\upsilon(p))$, and $\forall o_{1},o_{2}\in N\cup E,o_{1}\neq o_{2}\Rightarrow\sigma(o_{1})\cap\sigma(o_{2})=\varnothing$ and $\bigcup_{o\in N\cup E}\sigma(o)=P$.
Design decisions.
The constraints we have imposed in our design of MPGs are very minimalistic, ensuring only that all properties are uniquely assigned to edges and nodes, and that reification is well-founded. We have deliberately avoided imposing unwarranted constraints. For instance, there is no constraint preventing a node from having an outgoing edge to a subgraph containing one of the node’s own properties. This flexibility allows for representing scenarios such as “Mary legally changed her first name” or “Mary audits all names in the database (including her own)”. Similarly, there is no constraint requiring that for each edge in $\rho(n)$, the source and target of the edge must also be in $\rho(n)$. This permits representations like “Mary is confident that Bob edited a book, but not confident about which book”, where only the node representing Bob and Bob’s outgoing edge labeled ’edited’ are included, without the target of this edge. Of course, such additional constraints can be added (or existing constraints can be removed) as appropriate to the specific application domain.
It is worth noting that MPGs support meta-properties, i.e., properties on properties. For instance, we can represent “John’s birthdate was entered on 23 March 2020” since $\rho$ can consist of just a singleton containing a property. For simplicity, we have omitted special treatment of meta-properties in our main presentation.
We next formally define the graph substructure associated with a node.
Definition 4.2
(Sub-Structure) Let $G=(N,E,P,L,\lambda,\mu,\sigma,\upsilon,\eta,\rho)$ be a meta-property graph and $n\in N$. The sub-structure of $G$ induced by $n$ is $G_{n}=(N_{n},E_{n},P_{n},L_{n},\lambda_{n},\mu_{n},\sigma_{n},\upsilon_{n},{\eta}_{n},\rho_{n})$ where:
  • $N_{n}=N\cap\rho(n)$, $E_{n}=E\cap\rho(n)$, $L_{n}=L\cap\rho(n)$, $P_{n}=P\cap\rho(n)$
  • $\upsilon_{n}=\upsilon|_{P_{n}}$, $\mu_{n}=\mu|_{L_{n}}$ (domain restrictions)
  • $\lambda_{n}(o)=\lambda(o)$ for $o\in N_{n}\cup E_{n}$ if $\lambda(o)\in\rho(n)$, undefined otherwise
  • $\sigma_{n}(o)=\sigma(o)\cap\rho(n)$ for $o\in N_{n}\cup E_{n}$
  • ${E_{d}}_{n},{E_{u}}_{n}\subseteq E_{n}$ are directed and undirected edge subsets
  • ${\eta}_{n}=({\eta_{s}}_{n},{\eta_{t}}_{n},{\eta_{u}}_{n})$ where:
    • For $e\in{E_{d}}_{n}$:
        ${\eta_{s}}_{n}(e)=\eta_{s}(e)$ if $\eta_{s}(e)\in\rho(n)$, undefined otherwise
        ${\eta_{t}}_{n}(e)=\eta_{t}(e)$ if $\eta_{t}(e)\in\rho(n)$, undefined otherwise
    • ${\eta_{u}}_{n}(e)=\eta_{u}(e)\cap\rho(n)$ for $e\in{E_{u}}_{n}$
  • $\rho_{n}(n^{\prime})=\rho(n)\cap\rho(n^{\prime})$ for $n^{\prime}\in N_{n}$, satisfying the same well-foundedness property as in Definition 4.1.
Intuitively, the sub-structure $G_{n}$ can be thought of as a ”slice” or ”view” of the original meta-property graph $G$, focused on the node $n$ and its associated elements. This substructure includes only those nodes, edges, properties, and labels that are explicitly linked to $n$ through the reification function $\rho$. It maintains the original graph’s structure and relationships, but only for the elements within this subset. This allows us to zoom in on specific parts of the graph, making it easier to analyze and query complex, nested structures without losing the context of how they fit into the larger graph.
5 Meta-Property Graph Pattern Matching Language
We next formalize our Graph Pattern Matching Language for Metaproperty Graphs (MetaGPML) as an extension to the Graph Pattern Matching Language (GPML) [8, 12, 11]. GPML is the core element of both SQL/PGQ and GQL languages standardized by ISO. By extending GPML we ensure backward compatibility in the sense that every GPML query is a MetaGPML query.
5.1 Syntax
Let $\mathcal{X}$ be a countable infinite set of variables. We introduce descriptors $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ in MetaGPML, consistent with GPML[12]. For $x,y\in\mathcal{X},\ \ell\in\mathcal{L},\ \mathcal{k}\in\mathcal{K},$ and $\mathcal{v}\in\mathcal{V}$, we have
Descriptors: $\delta_{1}$ $:=x\mid:\ell\mid::\pi\mid x:\ell\mid:\ell::\pi\mid x::\pi\mid x:\ell::\pi$ (1) :?$y\mid x:?y\mid:?y::\pi\mid x:?y::\pi$ $\delta_{2}$ $:=x\mid:\ell\mid:?y\mid x:\ell\mid x:?y$ $\delta_{3}$ $:=x$
Here, $\pi$ is a pattern that describes different object patterns and some operations over them in a meta-property graph or sub-structure.
Patterns: $\begin{array}[]{llr}\pi:=&(\delta_{1})\ \mid\ ()\ \mid\ (\delta_{1}).\delta_{3}\ \mid\ ().\delta_{3}&\text{(node pattern)}\\ &\mid\ -[\delta_{2}]->\ \mid\ <-[\delta_{2}]-\ \mid-[\delta_{2}]-&\text{(edge pattern)}\\ &\mid\ -[]->\ \mid\ <-[]-\ \mid-[]-&\\ &\mid\ -[\delta_{2}].\delta_{3}->\ \mid\ <-[\delta_{2}].\delta_{3}-&\\ &\mid\ -[\delta_{2}].\delta_{3}-&\\ &\mid\ \{\delta_{3}\}\ \mid\ \{\}&\text{(property pattern)}\\ &\mid\ \ \mid\delta_{3}\mid\ \mid\ \mid\mid&\text{(label pattern)}\\ &\mid\ \pi\pi&\text{(concatenation)}\\ &\mid\ \pi+\pi&\text{(union)}\\ &\mid\ \pi\ \text{{WHERE}}\ \Phi&\text{(conditioning)}\\ \Pi:=&\pi\mid\Pi,\Pi&\text{(graph pattern)}\\ \end{array}$ (13)
We define $\Phi$ as the expression for conditions that can be used in a pattern after the WHERE keyword as follows, considering $p\in P$ and $c\in\mathcal{K}\cup\mathcal{V}\cup L$.
Expressions and $\begin{array}[]{llr}\mathcal{E}:=&\ x\mid\ x.\mathcal{k}\ \mid\ \mathcal{v}\ \mid\ \mathcal{k}\ \mid\ \ell&\hbox{\multirowsetup\text{(expressions)}}\\ &\ \mid\ \text{{KEY}}(p)\ \mid\ \text{{VAL}}(p)&\\ \Phi:=&\mathcal{E}=\mathcal{E}\ \mid\ \mathcal{E}<\mathcal{E}\ \mid\ x:\ell&\hbox{\multirowsetup\text{(conditions)}}\\ &\mid\text{{SUBSETEQ}}(x,\ y)\ \mid\ c\ \text{{ELEMENTOF}}\ y&\\ &\mid\ \Phi\ \text{{AND}}\ \Phi\ \mid\ \Phi\ \text{{OR}}\ \Phi\ \mid\ \text{{NOT}}\ \Phi&\end{array}$ (19)
Finally, based on (13), (19), we formulate the definition of clauses and queries.
Clauses and queries: $\begin{array}[]{ll}\mathcal{C}:=&\text{{MATCH}}\ \Pi\\ &\ \mid\ \text{{FILTER}}\ \Phi\\ \mathcal{Q}:=&\mathcal{C}\mathcal{Q}\\ &\mid\ \text{{RETURN}}\ \mathcal{E}_{1}\ \text{{AS}}\ x_{1},\ ...\ ,\mathcal{E}_{n}\ \text{{AS}}\ x_{n}\\ \end{array}$ (24)
The syntax of MetaGPML is designed to be intuitive and expressive, building upon familiar concepts from the GQL standard query language. Node and edge patterns use parentheses and brackets respectively, with labels prefixed by colons. The double-colon syntax (::) introduces subgraph patterns, enabling querying of reified structures. To distinguish them from constants, variables are prefixed with question marks when used to bind labels or properties. The language includes specialized predicates like ELEMENTOF and SUBSETEQ for working with sets of labels and properties.
5.2 Semantics
We next present the complete semantics of our language, as a clear unambiguous foundation for further study and implementation.
Definition 5.1
(Bindings) A binding $\beta:\mathcal{X}\rightarrow\mathbb{V}$ assigns variables $x\in\mathcal{X}$ to values $v\in\mathbb{V}$, where $\mathbb{V}=\mathcal{I}\cup\mathcal{L}\cup\mathcal{K}\cup\mathcal{V}$. We denote this as $(x_{1}\mapsto v_{1},...,x_{n}\mapsto v_{n})$, where $x_{1},...,x_{n}$ are the variables in the domain of $\beta$ $(\text{Dom}(\beta))$ and $v_{1},...,v_{n}$ are their corresponding values. We denote the empty binding as ()()\,( ).
Definition 5.2
(Compatibility of Bindings and their Join) Two bindings $\beta_{1},\beta_{2}$ are compatible, denoted by $\beta_{1}\sim\beta_{2}$, if they agree on their shared variables. Specifically, for every $x\in\text{Dom}(\beta_{1})\cap\text{Dom}(\beta_{2})$, it holds that $\beta_{1}(x)=\beta_{2}(x)$.
When $\beta_{1}\sim\beta_{2}$, their join $\beta_{1}\Join\beta_{2}$ is defined as follows: $\text{Dom}(\beta_{1}\Join\beta_{2})=\text{Dom}(\beta_{1})\cup\text{Dom}(\beta_{2})$. For any variable $x$, $(\beta_{1}\Join\beta_{2})(x)=\beta_{1}(x)$ if $x\in\text{Dom}(\beta_{1})\setminus\text{Dom}(\beta_{2})$, and $(\beta_{1}\Join\beta_{2})(x)=\beta_{2}(x)$ if $x\in\text{Dom}(\beta_{2})$.
A pattern matching semantic $\llbracket\pi\rrbracket_{G}$ in meta-property graph is a set of bindings $\beta$ that assign variables to values.
Node pattern matching: $\llbracket()\rrbracket_{G}=\{()\mid n\in N\}$ (25) $\llbracket(x)\rrbracket_{G}=\{(x\mapsto n)\mid n\in N\}$ (26) $\llbracket(x:\ell)\rrbracket_{G}=\{(x\mapsto n)\mid n\in N,\ell\in\mu(\lambda(n))\}$ (27) $\llbracket(x:?y)\rrbracket_{G}=\{(x\mapsto n,y\mapsto l)\mid n\in N,l=\lambda(n)\}$ (28) $\llbracket(x).z\rrbracket_{G}=\{(x\mapsto n,z\mapsto p)\mid n\in N,p\in\sigma(n)\}$ (29)
Equation 25 matches any node $n\in N$. Equation 26 assigns a node to variable $x$. Equation 27 matches nodes with specific labels, where $\mu(\lambda(n))$ determines the node’s labels. Equation 28 assigns a node’s label set to variable $y$. Finally Equation 29 marches nodes with property $p$ assigned to variable $z$.
Meta-nodes are a key concept in MetaGPML, defined as nodes to which diverse elements are assigned based on the concept of sub-structure. This feature enables reification within the Metaproperty Graph, allowing for a more nuanced representation of complex relationships and structures.
Equation 30 demonstrates the matching of nodes with labels to patterns within their associated sub-structure:
Meta-node pattern matching: $\llbracket(x:\ell::\pi)\rrbracket_{G}=\llbracket(x:\ell)\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (30)
The meta-node semantic definition involves a full join operation between two patterns divided by :::absent:::: : in the node pattern descriptor. Here, $\llbracket(x:\ell)\rrbracket_{G}$ represents a simple node pattern matching, while $\llbracket\pi\rrbracket_{G_{n}}$ is a pattern matching within the $G_{n}$ sub-structure. We define the semantics of edges in our model, supporting both directed and undirected edges.
Edge pattern matching: $\llbracket-[]->\rrbracket_{G}$ =$=\{()\mid e\in E_{d}\}$ (31) $\llbracket-[x]->\rrbracket_{G}$ =$=\{(x\mapsto e)\mid e\in E_{d}\}$ (32) $\llbracket-[:\ell]->\rrbracket_{G}$ =$=\{()\mid e\in E_{d},\ell\in\mu(\lambda(e))\}$ (33) $\llbracket-[x:\ell]-\rrbracket_{G}$ =$=\{(x\mapsto e)\mid e\in E_{u},\ell\in\mu(\lambda(e))\}$ (34) $\llbracket-[x:?y]->\rrbracket_{G}$ =$=\{(x\mapsto e,y\mapsto l)\mid e\in E_{d},l=\lambda(e)\}$ (35) $\llbracket-[x].z->\rrbracket_{G}$ =$=\{(x\mapsto e,z\mapsto p)\mid e\in E_{d},p\in\sigma(e)\}$ (36)
Equations 31-34 match various edge patterns, including directed and undirected edges with or without labels. Equation 35 allows assigning the edge’s label set to variable $y$, while 36 enables assigning the edge’s properties to variable $z$.
The meta-property graph model enables defining specific semantics for property and label objects, enhancing query capabilities on these elements.
Property and label objects: $\llbracket\{x\}\rrbracket_{G}$ =$=\{(x\mapsto p)\mid p\in P\}$ (37) $\llbracket|x|\rrbracket_{G}$ =$=\{(x\mapsto l)\mid l\in L\}$ (38)
Pattern concatenation, union, and conditioning: $\llbracket\pi_{1}\pi_{2}\rrbracket_{G}$ =$=\{\beta_{1}\Join\beta_{2}\mid\beta_{i}\in\llbracket\pi_{i}\rrbracket_{G},i=1,2\text{ and }\beta_{1}\sim\beta_{2}\}$ (39) $\llbracket\pi_{1}+\pi_{2}\rrbracket_{G}$ =$=\{\beta\cup\beta^{\prime}\mid\beta\in\llbracket\pi_{1}\rrbracket_{G}\cup\llbracket\pi_{2}\rrbracket_{G}\}$ (40) $\llbracket\pi\ \text{{WHERE}}\ \Phi\rrbracket_{G}$ =$=\{\beta\in\llbracket\pi\rrbracket_{G}\mid\llbracket\Phi\rrbracket^{\beta}_{G}=True\}$ (41)
In (40), $\beta^{\prime}$ maps any variable in ${\llbracket\pi_{1}+\pi_{2}\rrbracket}_{G}$ not in $\beta$’s domain to $Null$.
Graph patterns: $\llbracket\Pi_{1},\Pi_{2}\rrbracket_{G}=\{\beta_{1}\Join\beta_{2}\mid\beta_{i}\in\llbracket\Pi_{i}\rrbracket_{G},i=1,2\text{ and }\beta_{1}\sim\beta_{2}\}$ (42)
The semantics $\llbracket\mathcal{E}\rrbracket^{\beta}_{G}$ of an expression $\mathcal{E}$ is computed with respect to binding $\beta$ over graph $G$. For a variable $x$ in $\beta$’s domain, $\llbracket x\rrbracket^{\beta}_{G}=\beta(x)$.
Node and property values: $\llbracket x.\mathcal{k}\rrbracket^{\beta}_{G}=\begin{cases}Val(p)&\text{if }\beta(x)\in N\cup E,p\in\sigma(\beta(x)),Key(p)=\mathcal{k}\\ Null&\text{otherwise}\end{cases}$ (43)
Property key and value operations: $\llbracket\text{{KEY}}(p)\rrbracket^{\beta}_{G}$ $=Key(p)\text{ for }p\in P$ (44) $\llbracket\text{{VAL}}(p)\rrbracket^{\beta}_{G}$ $=Val(p)\text{ for }p\in P$ (45)
The semantics $\llbracket\Phi\rrbracket^{\beta}_{G}$ of a condition $\Phi$ is an element in $\{True,False,Null\}$, computed with respect to $\beta$ as follows.
Conditions: $\llbracket\mathcal{E}_{1}=\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ =$=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (46) $\llbracket\mathcal{E}_{1}<\mathcal{E}_{2}\rrbracket^{\beta}_{G}$ =$=\begin{cases}Null&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}=Null\text{ or }\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}=Null\\ True&\text{if }\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G}<\llbracket\mathcal{E}_{2}\rrbracket^{\beta}_{G}\\ False&\text{otherwise}\end{cases}$ (47) $\llbracket x:\ell\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&\text{if }\llbracket x\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\in\mu(\lambda(\llbracket x\rrbracket^{\beta}_{G}))\\ False&\text{if }\llbracket x\rrbracket^{\beta}_{G}\in N\cup E\text{ and }\ell\notin\mu(\lambda(\llbracket x\rrbracket^{\beta}_{G}))\end{cases}$ (48)
Logical operations: $\llbracket\Phi_{1}\ \text{{AND}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\wedge\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (49) $\llbracket\Phi_{1}\ \text{{OR}}\ \Phi_{2}\rrbracket^{\beta}_{G}$ $=\llbracket\Phi_{1}\rrbracket^{\beta}_{G}\vee\llbracket\Phi_{2}\rrbracket^{\beta}_{G}$ (50) $\llbracket\text{{NOT}}\ \Phi\rrbracket^{\beta}_{G}$ =$=\neg\llbracket\Phi\rrbracket^{\beta}_{G}$ (51)
Set operations, where $c\in\mathcal{K}\cup\mathcal{V}\cup\mathcal{L}$: $\llbracket c\ \text{{ELEMENTOF}}\ y\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&c\in\mu(\llbracket y\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (52) $\llbracket\text{{SUBSETEQ}}(x,y)\rrbracket^{\beta}_{G}$ =$=\begin{cases}True&\mu(\llbracket x\rrbracket^{\beta}_{G})\subseteq\mu(\llbracket y\rrbracket^{\beta}_{G})\\ False&\text{otherwise}\end{cases}$ (53)
Definition 5.3
MetaGPML evaluates queries and clauses using working tables, ensuring consistency with GQL standards. A table T comprises bindings with shared domains.
Finally, we define the semantics of clause $\mathcal{C}$ and query $\mathcal{Q}$ as functions operating on table T for graph $G$.
Clauses and queries: $\llbracket\text{{MATCH}}\ \Pi\rrbracket_{G}(T)$ $=\bigcup_{\beta\in T}\{\beta\Join\beta^{\prime}\mid\beta^{\prime}\in\llbracket\Pi\rrbracket_{G},\beta\sim\beta^{\prime}\}$ (54) $\llbracket\text{{FILTER}}\ \Phi\rrbracket_{G}(T)$ =$=\{\beta\in T\mid\llbracket\Phi\rrbracket^{\beta}_{G}=True\}$ (55) $\llbracket\mathcal{C}\ \mathcal{Q}\rrbracket_{G}(T)$ $=\llbracket\mathcal{Q}\rrbracket_{G}(\llbracket\mathcal{C}\rrbracket_{G}(T))$ (56) $,\displaystyle\llbracket\text{{RETURN}}\ \mathcal{E}_{1}\ \text{{AS}}\ x_{1},$$\ldots,\mathcal{E}_{n}\ \text{{AS}}\ x_{n}\rrbracket_{G}(T)=$ $\bigcup_{\beta\in T}\{(x_{1}\mapsto\llbracket\mathcal{E}_{1}\rrbracket^{\beta}_{G},\ldots,x_{n}\mapsto\llbracket\mathcal{E}_{n}\rrbracket^{\beta}_{G})\}$ (57)
The semantics of MetaGPML build upon and extend those of GPML, providing a robust framework for querying meta-property graphs. Key features include the ability to match nodes and edges with specific labels or properties, handle meta-nodes that encapsulate subgraphs, and perform operations on property and label objects. The language supports pattern concatenation, union, and conditioning, as well as set operations like ELEMENTOF and SUBSETEQ. Clauses such as MATCH and FILTER, along with the RETURN statement, allow for complex query construction. These foundations enable powerful and flexible querying capabilities, facilitating advanced analytics and metadata management in PG databases.
6 Next steps towards the MPG vision
In this paper we have highlighted the critical need to break down the barriers between data and metadata in property graph management. As first concrete steps towards this vision, we introduced a fully specified data model and query language for meta-property graphs, enabling the seamless modeling and interoperation of data and metadata. Our proposals are fully backwards compatible with the ISO standards for property graphs.
While this work lays the foundation for flexible management of property graphs for the requirements of contemporary applications, further research is needed to fully realize this vision. Next steps include investigating: (1) physical representations and indexing strategies for MPGs; (2) query evaluation and optimization solutions for the novel capabilities of MetaGPML; (3) extending the GQL and SQL/PGQ languages with MetaGPML; (4) schema and constraint languages for MPG, building upon the PG-SCHEMA framework [2]; and, (5) educational and training resources for students and professionals in working with MPGs and MetaGPML.